September 6, 2012 FDC meeting
- Production Construction Tracking (Dave)
- Grounding scheme (Vlad)
- Cathode corrosion
- Recent test results: 
- Cathode foil coating
- Engineering (Bill)
- Electronics (Chris, Nick)
- Third package tests FDC E-log (Beni)
- TDR, pages 132-145 (Lubomir)
Participants: Eugene, Bill, Dave, Simon, Vlad, Beni, and Lubomir
- Dave: Need to put the cards on another two cathodes and we will be ready with the first three cells that will go in package #1. Chris put the resistors on one PCB set and Casey started today making the PCB ring for wire frame #28. It will be laminated beginning of next week and Anatoly will put and test the HV caps before the end of the next week.
- Eugene: in order to expedite the solenoid testing, the plan so far is that three of our techs will be taken from us for two months, October and November. Dave and Anatoly will still work on the FDC. In case we need somebody for a short time, like half a shift, we should discuss this with Tim and Tom in advance.
- Just for the record: starting about Sunday evening or Monday morning (09/02-03) the humidity in the clean room was above 60%. On Tuesday it was 63% all day. A faulty fan was found that was replaced on Wednesday before lunch; in the afternoon the humidity returned back to normal for the season, 47-49%. In the clean room we have only the fourth package with four cell assembled being flushed with Nitrogen.
- Vlad updated the grounding scheme including the ~100 additional grounds, that is linked above.
- Another test with Viton+Apiezon was done in the vacuum chamber, this time using a spacer ring to have the right pressure on the O-ring applied with a lead brick. This time the lamps were kept on continuously over the holiday weekend; one of them burned on Monday. The copper changed color (picture attached). Unlike the previous test (with higher pressure) we could see only tiny holes through the copper using microscope and it's not not clear if they are a result of mechanical or chemical reaction.
- Olga analyzed the sample from the previous test, Viton+Apiezon at high pressure, and this time there's no sulfur visible, except at one place outside of the O-ring and Apiezon area, most likely pollution.
- We started several experiments coating the cathode foil below the O-ring:
- Using Kapton film glued with Hysol: doesn't work due to many bubbles and wrinkles.
- Coating only with Hysol: we used a "bad" cathode from package 1 and on the next day the Hysol spread away from the O-ring place that had some residue from the EPDM. Also, the coated surface was not very even. Trying to fix these problems we applied one more coating. We will also try this on a cathode that has not been damaged.
- Using Kapton film sitting on Apiezon (Anatoly's proposal); not real gluing but should work. The problem is that it will be very messy and Apiezon may accidentally go on the wires.
- Vlad proposed to study the rate of the chemical reaction (R) as a function of the temperature: R2=R1*Q^(T2-T1)/10degC. Q is between ~2 to 4 and can be measured doing our tests with the vacuum chamber at two different temperatures. The main question is how to measure the rate: we will measure the resistance and see how it changes with the time. The experiments should be done in normal conditions (other than the temperature) so that we can extrapolate the results.
- The tools (strong-backs) to move the cathodes work great using new and increased number (from 6 to 11) of suction cups.
Third package tests
- Beni made progress working with Cody in reading the fADC125 with the Linux ROC; time out is still there but can recover.
- The noise in the second cell was reduced by installing additional ground there. It turned out it was supposed to be there, just we missed to install it. We still can't use one negative HV sector and have two missing wire signals, because of the cards or chamber itself.
- Beni: gas system discussions in the control group: can save money installing less low-pressure sensors by multiplexing using manifolds.